OpenStudy (anonymous):

differentiate y=sin^-1[(5sinx + 4cosx)/√41]. i was stuck for 10 minutes on just this question in my test. it was so annoying!

hartnn (hartnn):

we can simplify this [(5sinx + 4cosx)/√41].

hartnn (hartnn):

know the formula for sin (A+B) ?

OpenStudy (anonymous):

dont remember it right now...

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how can you use sin(a+b) here?

OpenStudy (anonymous):

k, the formula is sin(a+b)=sin a . cos b + cos a . sin b

hartnn (hartnn):

like this [(5sinx + 4cosx)/√41] 5/√41 sin x + 4/√41 cos x let 5/√41 be cos t then 4/√41 = sin t so we have 5/√41 sin x + 4/√41 cos x = cos t sin x + sin t cos x = (sin (x+t)) got this ??

OpenStudy (anonymous):

how is 4/√41 = sin t ? is that part of the assumption ?

hartnn (hartnn):

where t is some constant

hartnn (hartnn):

assumption is " 5/√41 be cos t" from this can u find sin t ?

OpenStudy (anonymous):

uh, i dont see how you can get sin t from that.. do you use the identity sin^2+cos^2 ? (i hope im not sounding dumb)

hartnn (hartnn):

thats correct. use that identity you'll actually get sin t = 4/√41 that was the point of taking cos t as 5/√41 (and not just cos t = 5 or something)

hartnn (hartnn):

sin^2 t+cos^2 t =1

OpenStudy (anonymous):

okkkK! thanx a ton!!! but will i always be able to find sin t as a constant already present in the question? in these kind of sums?

hartnn (hartnn):

depends on the question actually, most of the times yes.

OpenStudy (anonymous):

and in the end when i have to resubstitute t, will it be like this cos t = 5/sqrt41, hence, t = cos^-1(5/sqrt41) ?

hartnn (hartnn):

since there is sqrt 41 term, i would say you can avoid it and simplify it further by finding tan t tan t = 4/5 right ? so you can put t= tan inverse (4/5) anyways, in this case you won't even need to re-substitute

hartnn (hartnn):

t = cos^-1(5/sqrt41) is also correct though

OpenStudy (anonymous):

At the end of the simplification part, it looks something like this y = sin^1. sin (x+t) so the sin^1 and sin get cancelled, and the eqn is y= x + t and when i differentiate it, i dont need to resubstitute t?

OpenStudy (anonymous):

its a constant

OpenStudy (anonymous):

so dy/dx will be just 1?

hartnn (hartnn):

correct 't' is some constant, not dependng on x so no need to re-substitute yes. dy/dx = 1 is correct

OpenStudy (anonymous):

ok! wow i dont believe the final answer could be so simple.. thanks a lot!

OpenStudy (primeralph):

@hartnn You're very sure right?

OpenStudy (anonymous):

but the sin^1 and sin do cancel each other out right?

hartnn (hartnn):

yes they do cancel. why @primeralph ? did you notice any error ?

OpenStudy (primeralph):

@hartnn Not really; I didn't go over it at all. I just wanted to know if I should.

hartnn (hartnn):

don't bother, its correct.

hartnn (hartnn):

@Stew.a.r.t. since you're new here, \(\large \text {Welcome to OpenStudy!!} \ddot \smile \) if you have any doubts browsing this site, you can ask me, or to an ambassador (with an encircled A before their name). Have fun learning with Open Study! :)

OpenStudy (anonymous):

k thnx